OpenStudy (anonymous):

Jack needs to earn a C. His test scores are 67, 72, 81, and 75. His final exam will count as three test scores. To earn a C, Jack's average must lie between 70 and 79(inclusive). What test score range can Jack earn on his final exam to receive a C?

OpenStudy (anonymous):

Let x = the required test score on the exam : Exam is worth 3 scores (3x), therefore we will be dealing with total of 7 test scores : (3x+67+72+81+75)/7= 70, his minimum for a C average (3x+295)/7= 70 Multiply both sides by 7 3x + 295 = 7(70) 3x + 295 = 490 3x = 490 - 295 3x = 195 x = 195/3 x = 65 min score on the exam for have a 70 average : Do the same with to find test score for 79 average (3x+67+72+81+75)/7 = 79, his minimum for a C average (3x+295)/7= 79 Multiply both sides by 7 3x + 295 = 7(79) 3x + 295 = 553 3x = 553 - 295 3x = 258 x = 258/3 x = 86 score on the exam for have a 79 average : The range of test scores would be from 65 to 86 for a 70 to 79 average .

OpenStudy (anonymous):

thanks.